OpenStudy (anonymous):

find the zeros of the function y=(x+2)(x-6)(x-7)

OpenStudy (anonymous):

There are three zeros, I'll give you the first one. x=-2, Can you find the rest?

OpenStudy (anonymous):

I dont know how to even do this

OpenStudy (anonymous):

okay. you pretty much have to solve for x, let me show you how.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Set the following equation equal to zero: y=(x+2)(x-6)(x-7), so (x+2)(x-6)(x-7)=0 so (x+2)=0 (x-6)=0 (x-7)=0 now solve each one for x. for the first one, x+2=0, subtract 2 from both sides, x=-2 The same goes for the rest. Let me know what you get.

OpenStudy (anonymous):

6 and 7

OpenStudy (anonymous):

Yes, x=6, 7, and -2

OpenStudy (anonymous):

That's really it.

OpenStudy (anonymous):

do you know how to write a polynomial in factored form

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

can you help me with x^3-2x^2-24x and writing it in factored form

OpenStudy (anonymous):

Sure thing. So first, start factoring out the x.

OpenStudy (anonymous):

x(x-6)(x+4)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

From x^3-2x^2-24x, after you factor out an x, you get x^2-2x-24.

OpenStudy (anonymous):

so, it looks something like x(x^2-2x-24). Then you factor the inside. In order to factor this, find what number you have to multiply to get 24( the last number), and what numbers you must add to get the inside number(-2)

OpenStudy (anonymous):

So if you multiply -6 by 4, you get -24, also if you add these two together you get -2

OpenStudy (anonymous):

Does that kinda help?

OpenStudy (anonymous):

Yeah thnaks